OpenStudy (emilyjones284):

medal and fan (: I just want to check my answer... A pizza restaurant offers nine choices of vegetarian toppings for pizza. In how many ways can you choose four toppings? is 15,120 correct?

jimthompson5910 (jim_thompson5910):

unfortunately that is incorrect how did you get that answer?

OpenStudy (emilyjones284):

i did 9! divided by 4!

jimthompson5910 (jim_thompson5910):

i see, you would use the formula n C r = (n!)/(r!*(n-r)!) where in this case n = 9 r = 4

jimthompson5910 (jim_thompson5910):

so... n C r = (n!)/(r!*(n-r)!) 9 C 4 = (9!)/(4!*(9-4)!) 9 C 4 = ??

OpenStudy (emilyjones284):

ohhhh so its 126?

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (emilyjones284):

thanks you i appreciate it! (:

jimthompson5910 (jim_thompson5910):

sure thign

jimthompson5910 (jim_thompson5910):

thing*

OpenStudy (emilyjones284):

thank* lol

jimthompson5910 (jim_thompson5910):

lol its ok, we both got typos...

OpenStudy (emilyjones284):

lol yeah now i dont feel as bad :P